OpenStudy (anonymous):

Eddie is 8 years older than his brother. The sum of their ages is 38, how old are they now?

jhonyy9 (jhonyy9):

do you have any idea to solve it ?

OpenStudy (anonymous):

All i did was 38 minus 8 but i know thats not correct. So no, i dont have any idea.

OpenStudy (ivancsc1996):

x+y=38 x-y=8 Solve the system.

OpenStudy (anonymous):

Oh so my original answer was right?

OpenStudy (ivancsc1996):

No, if you solve the system, one brother has 15 years and the other one 23

jhonyy9 (jhonyy9):

let brother age X than Eddie will be x+8 the sum of ages = 38 this mean that x+x+8=38 2x=30 x=15 x+8=15+8=23 Eddie age = 15 brother age =23 23+15=38

jhonyy9 (jhonyy9):

sorry inversed Eddie age is 23 and brither 15

OpenStudy (anonymous):

Ohh thanks :) i really had a long hard time trying to figure it out

jhonyy9 (jhonyy9):

brother 15

jhonyy9 (jhonyy9):

hope you understand it now sure right

OpenStudy (anonymous):

I sure do :) thanks so much!

jhonyy9 (jhonyy9):

yw